OpenStudy (anonymous):

Add. (3x^2 + 4x + 7) + (x^2 – 6x – 5) Answer 4x^2 – 10x + 2 3x^2 – 2x + 2 4x^x2 – 2x + 2 3x^2 + 2x – 2

OpenStudy (eyust707):

none of those

OpenStudy (anonymous):

(3x^2 + 4x + 7) + (x^2 – 6x – 5) Adding them term by term. 3x^2 +x ^2 = 4 x^2 4x-6x= -2x 7-5= 2 Putting it back tgt, 4x^2-2x+2

OpenStudy (anonymous):

The answer is C. U made a mistake typing one of the options.